Your answer is partially correct. Try again. In a
Answer #1

given that

L = 68 mH

C = 3.7 uF

we know that

capacitor will be fully charged first time when

t = T/4

where T = time period of oscillation

we know that

T = 2*pi / w

w = 1/sqrt (L*C)

so, t = pi*sqrt (L*C) / 2

t = 3.14 * sqrt (68*10^(-3)*3.7*10^(-6)) / 2

t = 7.86*10^(-4) s

answer